OpenStudy (itrymath):

The following function defines a recursive sequence. f(0) = -5 f(1) = 20 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? (4 points) -5, -20, -65, -200, … -5, 20, -92, 372, … -5, -24, -92, -372, … -5, 20, -65, 200, …

OpenStudy (itrymath):

@DanJS @zepdrix

OpenStudy (itrymath):

@3mar

zepdrix (zepdrix):

\[\large\rm f(\color{orangered}{n})=-4\cdot f(\color{orangered}{n}-1)-3\cdot f(\color{orangered}{n}-2)\] So umm.. you should be able to eliminate one of your options immediately. If our first term is -5, and the next is 20, you can ignore the list that doesn't start with -5 and 20.

zepdrix (zepdrix):

Plug in n=2 to find the third term of your sequence,\[\large\rm f(\color{orangered}{2})=-4\cdot f(\color{orangered}{2}-1)-3\cdot f(\color{orangered}{2}-2)\]\[\large\rm f(\color{orangered}{2})=-4\cdot f(1)-3\cdot f(0)\]

zepdrix (zepdrix):

They gave you the values for f(0) and f(1). So plug them in and simplify.

OpenStudy (itrymath):

whaaa

OpenStudy (itrymath):

this is so confusing

OpenStudy (3mar):

This is so easy! The first term is -5 The second term is 20 Any choice does not match these two criteria is excluded! Is it clear so far?

OpenStudy (itrymath):

what other choices are there?

OpenStudy (3mar):

-5, -20, -65, -200..........× -5, 20, -92, 372.............may be -5, -24, -92, -372...........× -5, 20, -65, 200.............may be

OpenStudy (itrymath):

oh okay so we can eliminate C

OpenStudy (itrymath):

and A

OpenStudy (3mar):

Congratulations!

OpenStudy (3mar):

@zepdrix Please proceed!

OpenStudy (3mar):

Where are you guys?

OpenStudy (itrymath):

oh sorry im back

OpenStudy (3mar):

So Are you satisfied?

OpenStudy (itrymath):

yes

OpenStudy (itrymath):

he wont come back i think so you will have to do next steps

OpenStudy (itrymath):

@3mar can we work on this together quickly

OpenStudy (3mar):

Pleasure is mine!

OpenStudy (3mar):

Yes, can you?

OpenStudy (itrymath):

yes wait 2 min

OpenStudy (itrymath):

please thanks

OpenStudy (3mar):

As @zepdrix said and mentioned: \[\large\rm f(\color{orangered}{2})=-4\cdot f(\color{orangered}{2}-1)-3\cdot f(\color{orangered}{2}-2)\] \[\large\rm f(\color{orangered}{2})=-4\cdot f(1)-3\cdot f(0)\]

OpenStudy (3mar):

Ca you find out the third term!?

OpenStudy (itrymath):

@3mar sooo sorry